OpenStudy (anonymous):

A 1-dimensional elastic collision occurs between two masses m1 = 3 kg and m2 = 5 kg. If the speeds after the collision are v1f = +3 m/s and v2f = -2 m/s, what were their speeds v1i and v2i before the collision?

OpenStudy (anonymous):

I was setting up two equations. One with momentum and the other with kinetic energy

OpenStudy (anonymous):

momentum 3v1+5v2=3(3)-5(2) energy .5(3)v1^2+.5(5)v2^2=.5(3)3^2+.5(5)2^2